Solve: e^(x)sinydx + (e^x +1)cosydy=0
1
Expert's answer
2020-09-02T19:06:44-0400

We separate the variables and receive the following:


"\frac{cos\,y}{sin\,y}dy+\frac{e^x}{e^x+1}dx=0"

After integration we receive "ln(sin\,y)+ln(e^x+1)+c=0,\,\,c\in\mathbb{R}"

From this we receive that "ln(sin\,y(e^x+1)e^c)=0" . Thus, "sin\,y(e^x+1)e^c=1"

The latter implies that "sin\,y=\frac{1}{(e^x+1)e^c}" . From the latter we find that "y=(-1)^narcsin\left(\frac{1}{(e^x+1)e^c}\right)+\pi n."

Answer: "y=(-1)^narcsin\left(\frac{1}{(e^x+1)e^c}\right)+\pi n" , "c\in\mathbb{R}"